OpenStudy (anonymous):

If 3/4 of a cup of food gives you your daily value of potassium, then what fraction of your daily value of potassium is in 1/2 cup of the food? Is the answer 1/3?

OpenStudy (anonymous):

try looking at it through proportions.... if 3/4 cups gives you 100% of your daily allowance, then 1/2 cup gives you x: (3/4)/100 = (1/2)/x solving this should give you 200/3 percent or 66 2/3 % = 2/3.

OpenStudy (anonymous):

The only problem is that I cannot use a calculator during the exam.

OpenStudy (anonymous):

you don't need a calculator for this problem... those are "easy" numbers...